OpenStudy (anonymous):

Combine like terms. What is a simpler form of the expression? -3(-4y + 3) + 7y A. 19y - 9 B. 10y C. -19y + 3 D. -19y - 9

OpenStudy (anonymous):

i think it's D

OpenStudy (cwrw238):

first use the distributive property and add like terms

OpenStudy (cwrw238):

-3 (-4y + 3) = -3*-4y + -3*3 = 12y - 9 remember: - multiplied by - = +

OpenStudy (cwrw238):

not D